# Enhanced Continual Learning of Vision-Language Models with Model Fusion

### v1

@@ 3 Experiments
Table 1: Comparison with SOTA methods on MTIL benchmark in terms of “Transfer”, “Average.”, and “Last” scores (%). We label the best methods on average of all datasets with bold styles. The results of more baselines can be found in Appendix I.
@@
The detailed experiment setting is shown in Appendix B. We report two results of ConDU. ConDU(FT) denotes ConDU with full fine-tuning, and ConDU(LoRA) denotes ConDU with LoRA-based fine-tuning. The experimental results on task-agnostic MTIL is shown in I.
Table 2: Comparison with SOTA methods on few-shot MTIL benchmark in terms of “Transfer”, “Average.”, and “Last” scores (%). We label the best methods on average of all datasets with bold styles. The results of more baselines can be found in Appendix I.
@@ Appendix D Related Work
Appendix E Introduction of LoRA
Below we revisit LoRA, in which $g$ is the module that PEFT attached to, $\mathbf{e}$ and $\mathbf{h}$ are input and output of the original $g$ and $\mathbf{h}^{\prime}$ is output of $g$ attached with PEFT.
LoRA Hu et al. (2021) assumes the change of parameters is in a low-rank space when tuning the pre-trained model on a downstream task. For a linear layer with weight $\mathbf{W}\in\mathbb{R}^{d\times d^{\prime}}$, the weight updates $\Delta\mathbf{W}$ can be decomposed into the multiplication of two small matrices:
$\Delta\mathbf{W}=\mathbf{W}_{down}\mathbf{W}_{up},$
where $\mathbf{W}_{down}\in\mathbb{R}^{d\times r}$ and $\mathbf{W}_{up}\in\mathbb{R}^{r\times d^{\prime}}$. For the convolution layer, the updates can be reshaped into the kernel shape. Finally, LoRA modifies the forward pass of the adapted layer into the following form:
$\textbf{h}^{\prime}=\textbf{h}+\textbf{e}\ast(\mathbf{W}_{down}\mathbf{W}_{up}),$
where $\ast$ is matrix multiplication or convolution operation, the bias and reshape operation are omitted for conciseness. Since LoRA adapts the weight of $g$, the weight updates can be merged into $g$ to reduce the inference latency.
Appendix F Storage Analysis
@@
Appendix G Theoretical Analysis
G.1 Definitions
@@
G.2 Related Lemmas
Lemma G.1 (Sign Preservation of $\delta^{i}$).
@@ G.2 Related Lemmas
Lemma G.2 (Preservation of $L_{1}$ Norm).
@@
G.3 Main Theorems
Theorem G.3 (Convergence of Iteration).
@@ G.3 Main Theorems
Indeed, from the proof of Lemma G.1, we know that the number of zeros in $M^{i}$ does not decrease during iterations. Since no new delta models are introduced, the number of zeros in $M^{i}$ cannot increase either. Thus, the sign of each position remains fixed during iterations.
@@
Next, we consider $i_{2}$ and divide $\delta^{i_{2}}$ into two parts. Let $x_{t}$ denote the sum of absolute values in $\delta^{i_{2}}$ at positions where $M^{i_{1}}=1$ during the $t$-th iteration, and let $y_{t}$ represent the sum of absolute values at positions where $M^{i_{1}}=0$. By Lemma G.2, we know that $x_{t}+y_{t}=c$ for all $t$, where $c$ is a constant. During each iteration, the values in $y_{t}$ are incorporated into $\epsilon_{uni}$. Let $s$ denote the sum of absolute values in $\delta^{i_{1}}$ at positions where $M^{i_{2}}=1$, which is also a constant.Then, we have:
@@
Corollary G.4.
Under the same conditions as Theorem G.3, the following holds:
@@
Theorem G.5.
If an initial delta model $\delta^{1}$ is given, and during the $n$-th operation, a new delta model $\delta^{n}$ is added, and the current set of delta models $\{\delta^{i}(n)\mid i\in\{1,\dots,n+1\}\}$ undergoes one iteration, then under the same conditions as Theorem G.3, and assuming all $\delta^{i}$ are independent and identically distributed, we have:
@@
For (1): From Lemma G.1, if $M^{i}_{k}(j)$ changes, it must transition from 1 to 0. This implies that the sign of $\delta^{n}_{k}$ is different from that of $\delta^{i}_{k}$, and $\left\lvert\delta^{n}_{k}\right\rvert\geq\sum_{i=1}^{n-1}\left\lvert\delta^{i}_{k}(n-1)\right\rvert$. Under the assumption of independence and identical distribution, this probability approaches 0 as $n$ increases.
For (2): Based on the proof of Theorem G.3, if $\epsilon_{uni}$ changes its selected corresponding delta model at any position during an iteration, this change can only be caused by one of the two most recently added delta models $\delta^{n}$. There are only two possible cases:
@@
Corollary G.6.
Under the same conditions as Theorem G.5, the following holds:
@@
From Theorem G.3, after one iteration, all delta models gradually stabilize. From Theorem G.5, the impact of newly added delta models on the iteration process is negligible. Thus, Corollary G.4 also implies Corollary G.6. ∎
Appendix H Ablation Study of $K$ of Aggregating
@@
Appendix I More Results of Comparison with State-of-the-art Methods
@@
Appendix J Other Order of MTIL
@@
Appendix K $t$-SNE Visualization of Feature Space

### v2

@@ 3 Experiments
Table 1: Comparison with SOTA methods on MTIL benchmark in terms of “Transfer”, “Average.”, and “Last” scores (%). We label the best methods on average of all datasets with bold styles. The results of more baselines can be found in Appendix H.
@@
The detailed experiment setting is shown in Appendix B. We report two results of ConDU. ConDU(FT) denotes ConDU with full fine-tuning, and ConDU(LoRA) denotes ConDU with LoRA-based fine-tuning. The experimental results on task-agnostic MTIL is shown in H.
Table 2: Comparison with SOTA methods on few-shot MTIL benchmark in terms of “Transfer”, “Average.”, and “Last” scores (%). We label the best methods on average of all datasets with bold styles. The results of more baselines can be found in Appendix H.
@@ Appendix D Related Work
Appendix E Storage Analysis
@@
Appendix F Theoretical Analysis
F.1 Definitions
@@
F.2 Related Lemmas
Lemma F.1 (Sign Preservation of $\delta^{i}$).
@@ F.2 Related Lemmas
Lemma F.2 (Preservation of $L_{1}$ Norm).
@@
F.3 Main Theorems
Theorem F.3 (Convergence of Iteration).
@@ F.3 Main Theorems
Indeed, from the proof of Lemma F.1, we know that the number of zeros in $M^{i}$ does not decrease during iterations. Since no new delta models are introduced, the number of zeros in $M^{i}$ cannot increase either. Thus, the sign of each position remains fixed during iterations.
@@
Next, we consider $i_{2}$ and divide $\delta^{i_{2}}$ into two parts. Let $x_{t}$ denote the sum of absolute values in $\delta^{i_{2}}$ at positions where $M^{i_{1}}=1$ during the $t$-th iteration, and let $y_{t}$ represent the sum of absolute values at positions where $M^{i_{1}}=0$. By Lemma F.2, we know that $x_{t}+y_{t}=c$ for all $t$, where $c$ is a constant. During each iteration, the values in $y_{t}$ are incorporated into $\epsilon_{uni}$. Let $s$ denote the sum of absolute values in $\delta^{i_{1}}$ at positions where $M^{i_{2}}=1$, which is also a constant.Then, we have:
@@
Corollary F.4.
Under the same conditions as Theorem F.3, the following holds:
@@
Theorem F.5.
If an initial delta model $\delta^{1}$ is given, and during the $n$-th operation, a new delta model $\delta^{n}$ is added, and the current set of delta models $\{\delta^{i}(n)\mid i\in\{1,\dots,n+1\}\}$ undergoes one iteration, then under the same conditions as Theorem F.3, and assuming all $\delta^{i}$ are independent and identically distributed, we have:
@@
For (1): From Lemma F.1, if $M^{i}_{k}(j)$ changes, it must transition from 1 to 0. This implies that the sign of $\delta^{n}_{k}$ is different from that of $\delta^{i}_{k}$, and $\left\lvert\delta^{n}_{k}\right\rvert\geq\sum_{i=1}^{n-1}\left\lvert\delta^{i}_{k}(n-1)\right\rvert$. Under the assumption of independence and identical distribution, this probability approaches 0 as $n$ increases.
For (2): Based on the proof of Theorem F.3, if $\epsilon_{uni}$ changes its selected corresponding delta model at any position during an iteration, this change can only be caused by one of the two most recently added delta models $\delta^{n}$. There are only two possible cases:
@@
Corollary F.6.
Under the same conditions as Theorem F.5, the following holds:
@@
From Theorem F.3, after one iteration, all delta models gradually stabilize. From Theorem F.5, the impact of newly added delta models on the iteration process is negligible. Thus, Corollary F.4 also implies Corollary F.6. ∎
Appendix G Ablation Study of $K$ of Aggregating
@@
Appendix H More Results of Comparison with State-of-the-art Methods
@@
Appendix I Other Order of MTIL
@@
Appendix J $t$-SNE Visualization of Feature Space